OpenStudy (anonymous):

Diego's mother is twice as old as he is she is also as old as the sum of the ages of diego and both of his younger twin brothers. the twins are 11years old solve the equation 2d=d+11+11 to find he age of diego

OpenStudy (anonymous):

2d-d=d+11+11-d d=11+11 d=22 therefore, diego is 22yrs old